x2 + y2 = 8
x - y = 0
find the solutions to the system

Found 2 solutions by Alan3354, mananth:
Answer by Alan3354(69443)

x = y = ± 2

Answer by mananth(16949)

x2 + y2 = 8
x - y = 0
find the solutions to the system
x^2+y^2=8
x^2-2xy+y^2=8-2xy
(x-y)^2=8-2xy
substitute (x-y) =0
(0)^2=8-2xy
2xy=8
xy=4
x=4/y
...
x-y =0
4/y - y =0
(4-y^2)/y=0
4-y^2=0
y^2=4
take the square root
y= +/- 2
(2,-2) & (-2,2)
m.ananth@hotmail.ca